Can you assist me in solving this word problem?
An object's altitude in meters is given by the polynomial h + vt - 9.8t^2, where h is the height in meters from which the rocket is airborne. A pebble is shot upward from the top of a building 200 meters tall. If the initial speed is 30 meters per second, how high above the ground will the pebble be after 4 seconds? (Round results to the nearest tenth of a meter)

Answer by Alan3354(69443)

An object's altitude in meters is given by the polynomial h + vt - 9.8t^2, where h is the height in meters from which the rocket is airborne. A pebble is shot upward from the top of a building 200 meters tall. If the initial speed is 30 meters per second, how high above the ground will the pebble be after 4 seconds?
-------------
h(t) = 200 + 30t - 9.8t^2
h(4) = 200 + 120 - 9.8*16
h(4) = 163.2 meters
-----------------------
9.8 meters/sec/sec is the acceleration due to gravity
The term in the equation should be 1/2 that, or 4.9 m/sec/sec since
s = (at^2)/2, but that's the value given. Maybe this takes place on a different planet.